OpenStudy (anonymous):

if f(g(x))=x and f(x)=3x-4,find g(x)

OpenStudy (anonymous):

is it g(x)=x+4/3

OpenStudy (anonymous):

i am going to assume G is the inverse function of F(x) so to find the inverse function let f(x)= y so you get y=3x-4 to get the inverse function you want to make x the subject add 4 to both sides y+4=3x divide by 3 on both sides (y+4)/3=x now let x=g(x) let y=x (x+4)/3=g(x)

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

UR SMART

OpenStudy (anonymous):

haha i try lol